Progress


The decomposition reaction of NOBr is second order in NOBr, with a rate constant at 20°C of 25 M1 min 1. If the initial concentration of NOBr is 0.025 M, find


(a) the time at which the concentration will be 0.020 M.


1
Expert's answer
2022-01-30T07:58:26-0500

NOBr decomposition follows the equation 2NOBr = 2NO + Br2. Thus the rate equation for the second order reaction will be k2 = 1/t(1/C - 1/C0). Using this equation we can calculate the the time of the concentration at 0.020M


C = C0/(C0*k2*t + 1) = 0.025 M/(0.025 M*25 M-1 min-1*x + 1) = 50minutes